OpenStudy (anonymous):

which ordered pair is a solution to the equation y=2x+8 (2,1) (-2,-3) (0,8) (8,0)

OpenStudy (lucaz):

do you have options?

OpenStudy (phi):

test each choice. that means for (2,1) take the first number (the 2) and use that instead of x in your equation y = 2*x + 8 replace the x with 2: y = 2*2 + 8 y = 4+8 y=12 now compare the 12 you get with the second number in (2,1) are they the same number? NO. First choice does not work. Can you try the next choice ?

OpenStudy (anonymous):

0,8